the perimeter of a rectangle is 56 cm. The width is 6 cm less than the length. Find the width and the length.
Answer by rapaljer(4671)

Let x= length (because this is the last phrase in the first sentence!)
x-6 = width

2W + 2L = Perimeter
2(x-6) +2(x)= 56
2x-12+2x=56
4x-12=56

Add +12 to each side:
4x-12+12=56+12
4x=68
x=17 Length
x-6=11 Width

Check:
2*17+2*11
34+22
56 It checks!!

For more help with WORD PROBLEMS, in particular, Perimeter problems like this one, please see my website. Click on my tutor name "Rapaljer" anywhere in algebra.com, which takes you to the address of my website. From the Homepage of my website, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time." Choose "Basic Algebra", and look in "Chapter 1" for "Section 1.10 Word Problems." Here I have tried to give a non-traditional explanation designed for students who have trouble with math. My own students have unanimously told me that they understood MY explanations a LOT better than traditional textbooks! On this topic, I have quite a lot of the exercises solved IN COLOR on my "MATH IN LIVING COLOR" pages that accompany my explanations. Try it, and see if it works for you.

Then, if you like the website, try it for other topics as well, and recommend it for your friends and family members. It's all FREE, just like algebra.com!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus